Part A A particle travels along a straight line with a velocity of v (18 0.05s) m/s, where s is in meters. Determine the acceleration of the particle at s 10 m Express your answer to three significant figures and include the appropriate units. a-Value m/s Submit Request Answer
Answer #1

Velocity of particle is given by:

V = 18 - 0.05*s^2

We know that acceleration is given by:

a = dV/dt

Also velocity is given by:

V = ds/dt

So differentiating above equation w.r.t to t

V = 18 - 0.05*s^2

dV/dt = 0 - 2*0.05*s*(ds/dt)

a = -0.1*s*V

a = -0.1*s*(18 - 0.05*s^2)

a = 0.005*s^3 - 1.8*s

Now when s = 10 m

a = 0.005*10^3 - 1.8*10

a = -13.0 m/sec^2 = acceleration of the particle.

Please Upvote.